# Improving Multilingual Capabilities with Cultural and Local Knowledge in Large Language Models While Enhancing Native Performance

### v1

@@ Abstract
Large Language Models (LLMs) have shown remarkable capabilities, but their development has primarily focused on English and other high-resource languages, leaving many languages underserved. We present our latest Hindi-English bi-lingual LLM Mantra-14B with  3% average improvement in benchmark scores over both languages, outperforming models twice its size. Using a curated dataset composed of English and Hindi instruction data of 485K samples, we instruction tuned models such as Qwen-2.5-14B-Instruct and Phi-4 to improve performance over both English and Hindi. Our experiments encompassing seven different LLMs of varying parameter sizes and over 140 training attempts with varying English-Hindi training data ratios demonstrated that it is possible to significantly improve multilingual performance without compromising native performance. Further, our approach avoids resource-intensive techniques like vocabulary expansion or architectural modifications, thus keeping the model size small. Our results indicate that modest fine-tuning with culturally and locally informed data can bridge performance gaps without incurring significant computational overhead. We release our training code, datasets, and models under mit and apache licenses to aid further research towards under-represented and low-resource languages.
@@ 2 Related Works
Prior studies have attempted to address this disparity through various techniques, including vocabulary expansion/modification (Tejaswi et al., 2024) (Csaki et al., 2023) (Shi et al., 2024) (Balachandran, 2023), modifications in architecture (Llama-Nanda, 2024) like block expansion and the addition of extra layers to accommodate linguistic diversity, or continued pre-training followed by instruction tuning again (Mahdizadeh Sani et al., 2025) (Kuulmets et al., 2024) (Cui et al., 2023) (Vo et al., 2024) (Luukkonen et al., 2023) (Toraman, 2024). However, such methods often incur substantial computational costs and lead to increase in model sizes.
Some more closely relevant works would be models that are optimized for Hindi (Aryabumi et al., 2024), (Dang et al., 2024), (Üstün et al., 2024), (BigScienceWorkshop, 2023) and other mono-lingual and bi-lingual LLMs focused on Hindi (Llama-Nanda, 2024), (Gala et al., 2024), (BhabhaAI, 2024), (GenVRadmin, 2024), (Joshi et al., 2024).
@@ 3 Datasets
Despite the existence of datasets to cover several domains for Hindi (Khan et al., 2024), (Ramesh et al., 2022), we decided to experiment primarily with translated / reformatted datasets which do not prohibit usage for research/commercial purpose. This was done so that the same work can be implemented/extended to low-resource languages. For translation, we used GPT-4o-mini (OpenAI, 2024) through Microsoft Azure  [^1] https://azure.microsoft.com/en-us/products/ai-services/openai-service/ to translate few datasets and benchmarks from English to Hindi : Big-Bench-Hard (Suzgun et al., 2022), XNLI (Conneau et al., 2018), Xl-Sum (Hasan et al., 2021). Some of the benchmarks which already have Hindi subsets were used directly : Global MMLU (Singh et al., 2024a), IndicXNLI (Aggarwal et al., 2022).Some of the publicly available datasets containing cultural and localized general knowledge like Indian legal FAQ(Aditya2411, 2024), UPSC FAQ (prnv19, 2024), IndianTAX FAQ (msinankhan1, 2024), IndianMedicines , IndiaCuisines and IndiaTravel Guide (cyberblip, 2024) were used to generate instruction-response pairs from the tabular format data using GPT-4o-mini as a part of our dataset collection. These were first translated to the other language from the original language then manually verified by multiple annotators to ensure quality in both languages. We also used a few subsets from the Aya collection (Singh et al., 2024b) i.e the translation, simplification and summarization subsets. In total the collected dataset had 3.12M samples with nearly 50:50 ratio of English and Hindi data. Around 90K samples from these cover localized and cultural knowledge. Among the rest, some domains and tasks had higher proportion in the collection. We used randomly selected subsets from those datasets while maintaining equal language ratios. After filtering the training data, we had around 485K samples of which 20% are of localized domain and cultural knowledge, while the rest are of generic tasks like math, MCQs, reasoning, summarization, rephrasing and translation.
@@ 6 Dataset Distribution and Ordering
The performance of models from initial tests didn’t vary significantly with/without being trained on math data. The performance on Math subsets of MMLU as well remained similar on both languages with/without being trained on math samples. Since we would be training on a large number of samples, we decided to still use a considerable amount of math samples. A significant performance gap was observed over boolean benchmarks with a nearly 3% increase in English and 5% increase in Hindi. Hence, we decided to use a slightly higher amount of boolean questions’ samples in the final dataset. The language ratios for each domain in the final dataset were determined based on the initial training data ratios that gave the best results. The samples of the final dataset were sorted over input lengths in ascending order with a certain number of longest samples placed in the beginning. This number was set equal to the total effective batch size (i.e the product of batch size and gradient accumulation steps). The samples related to local and cultural knowledge were then placed such that they are evenly spread out in the dataset except the initial batch. More info on the dataset can be found in Appendix B. The training methods and details can be found in Appendix A.
@@ 7 End Evaluation
Apart from the benchmarks seen in Table 2 and Table 3, we perform evaluations over additional benchmarks like : MMLU-Pro (Wang et al., 2024), BigBench-Hard (Suzgun et al., 2022), MuSR (Sprague et al., 2024), GPQA (Rein et al., 2023), MATH-Hard (Hendrycks et al., 2021). We used open-llm-leaderboard  [^2] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard (Fourrier et al., 2024) for evaluation over some of the benchmarks through eval-harness framework(Gao et al., 2021). Table 7 demonstrates The performance of our models in comparison with the original models over several benchmarks. We did observe variations in the scores from open-llm-leaderboard and the corresponding benchmark scores which were self reported for the original models. We used the scores from the leaderboard for all models over those benchmarks for reproducibility a fair comparison. The evaluation methods used can be seen in Table 4.
@@
Figure 1: Model Performance (Avg. benchmark score) VS Model Size for English (left) and Hindi (right)
@@
Our Qwen Model (14b) 90.61 94.82 88.53 90.70 75.00 87.93 52.63 36.24 44.84 64.97 25.08
Mantra-14B 97.39 92.24 87.65 87.40 75.59 88.05 52.39 39.77 49.07 66.97 23.11
@@
Our Qwen Model (14b) 74.06 81.23 84.07 78.20 53.85 74.82
Mantra-14B 81.74 89.06 86.02 78.70 56.39 78.38
@@ 8 Comparisons
For additional comparisons, we compare the performance of our models with other Hindi bilingual LLMs and other open-source LLMs which are optimized for Hindi. Due to the large variations in number of parameters of our models and other comparable models, we compare average benchmark performance versus the model size in terms of VRAM requirement. The comparisons over English and Hindi benchmarks along side our Qwen and phi models can be seen in Table 5 and Table 6. The comparison of size of models to average of benchmarks scores can be seen in Figure 1. Over the benchmarks of higher difficulty, our models have consistently outperformed models over twice their size as seen in Table 5.
Benchmark Lang Qwen-2.5- Our Qwen Change Phi-4 Mantra-14B Change
@@
Benchmark Lang Qwen-2.5- Our Qwen Change Phi-4 Mantra-14B Change
@@ 8.1 Domain wise Performance change
The performance of our models compared to the original versions over MMLU-pro can be seen in Table 9. The type of questions the models faced through MMLU-Pro maybe of the same domain but were of different subdomains and task types compared to those in our datasets. For example, The CS benchmarks’ questions were MCQs about various areas of computer science while our training data over CS was solely from MBPP (Austin et al., 2021) which consists of a text input and a python code as an output. Further the only source of training data we used for economics consist of TAX filing FAQs over Indian context and primarily in Hindi. Hence such domains’ data usage was mentioned as N/A. The domains which had a performance boost in our models without being in training data had questions of the form of fill-mask or text completion which were similar to the training data taken from Winogrande-XL (Sakaguchi et al., 2021) and PIQA (Bisk et al., 2020) spanning several domains.
@@
Figure 2: Distribution of each model’s choices over MMLU-Pro
@@ 8.2 Model biases over choices
The observations from domain wise performance changes by Phi and Qwen were significantly different. The domains which were well represented in our training data had a significant boost on both languages of MMLU. Despite training on MCQs which consist of 2-4 options, similar results of improvement were seen over MMLU-Pro which has upto 10 options. On the other hand, Phi-4 had a higher performance boost over MMLU which has the same number of options as the samples in the training data, but the performance over MMLU-Pro dropped irrespective of domain. The distribution of choices made by each of our LLMs and the corresponding original implementation can be seen in Figure 2. The instruction tuning dataset we used had an equal distribution of each of the choices among MCQ samples. The original Qwen model overwhelmingly chose from the final two options while our model was able to generalize well despite not being trained on MCQs with 10 choices, with an inclination towards the 5th option. On the other hand the original phi-4 was able to perform better than its counterpart, but despite being fine-tuned with equal distribution of choices, the model displayed an inclination towards the first choice among the list of options. The extent of this bias varied between each domain significantly. More on this can be seen in Appendix D. As our models were fine-tuned from the original models’ instruct variants, the biases were assumed to have been carried forward. Our models were able to respond well with less biases in choices over the domains whose samples are present in large quantities in our training data. To further look into this, we tried to fine-tune the base version of qwen-2.5-14B rather than the instruct model to see the choices made on MMLU-Pro, while most of our dataset’s samples of MCQs were having 4-5 samples, it was reflected in the choices made as seen in Figure 3 which demonstrates the issue within the original model similar to previous works demonstrating sensitivity on models’ sensitivity to order of choices (Pezeshkpour and Hruschka, 2024). But a well balanced instruction tuning dataset can minimize this issue or an evaluation independent of order of choices (Zheng et al., 2023). A slight tilt from left to right in Figure 2 and Figure 3 can be expected as not all questions are accompanied by 10 options with a considerable amount having less.
Figure 3: our custom instruction tuned model’s choices over MMLU-Pro
@@ 9 Conclusion
We demonstrate that enhancing low-resource language capabilities in LLMs is possible through targeted fine-tuning rather than complex architectural changes. Our work shows that a 12-15B parameter LLM provides an effective balance between performance and accessibility, requiring just 30GB RAM in bf16 or 10GB of RAM in 4-bit quantization. The performance analysis reveals that our Phi-4 model excels in general-purpose tasks, while the Qwen model shows stronger adaptation to specific domains, as evidenced by the domain-wise performance changes in Table 9. Our approach of using primarily translated datasets, except for culturally specific knowledge, makes this method readily adaptable to other low-resource languages. To further push the research in low-resource languages, we release our training code, datasets, and models under commercially permissible licenses.
@@ 10 License
Our Qwen and Phi models are available through the same licenses as the models we used as a base i.e apache-2.0 and mit respectively. the models can be accessed here  [^3] Our Phi-4 model :https://huggingface.co/large-traversaal/Mantra-14B. The training datasets are publicly available here [^4] Datasets : 1024m/PHI-4-Hindi-Instruct-Data. Most datasets used for training the models have a copyleft license, with the rest having no license specified and are publicly available on huggingface.
@@ Appendix A Model Replication
The hyper-parameters used for training can be seen below in Table 10. The inital training attempts using a portion of the data (i.e 8% samples) were done on various different devices, the final models were trained on a single H200 SXM for 55,56,54 hours each using Qwen2.5-14B-Instruct, Phi-4, Qwen2.5-14B-base respectively.

### v2

@@ Abstract
Large Language Models (LLMs) have shown remarkable capabilities, but their development has primarily focused on English and other high-resource languages, leaving many languages underserved. We present our latest Hindi-English bi-lingual LLM with  3% average improvement in benchmark scores over both languages, outperforming models twice its size. Using a curated dataset composed of English and Hindi instruction data of 485K samples, we instruction tuned models such as Qwen-2.5-14B-Instruct and Phi-4 to improve performance over both English and Hindi. Our experiments encompassing seven different LLMs of varying parameter sizes and over 140 training attempts with varying English-Hindi training data ratios demonstrated that it is possible to significantly improve multilingual performance without compromising native performance. Further, our approach avoids resource-intensive techniques like vocabulary expansion or architectural modifications, thus keeping the model size small. Our results indicate that modest fine-tuning with culturally and locally informed data can bridge performance gaps without incurring significant computational overhead. We release our training code, datasets, and models under mit and apache licenses to aid further research towards under-represented and low-resource languages.
@@ 2 Related Works
Prior studies have attempted to address this disparity through various techniques, including vocabulary expansion/modification (Tejaswi et al., 2024) (Csaki et al., 2023) (Shi et al., 2024) (Balachandran, 2023), modifications in architecture like block expansion and the addition of extra layers to accommodate linguistic diversity (Llama-Nanda, 2024), or continued pre-training followed by instruction tuning again (Mahdizadeh Sani et al., 2025) (Kuulmets et al., 2024) (Cui et al., 2023) (Vo et al., 2024) (Luukkonen et al., 2023) (Kallappa et al., 2025) (Toraman, 2024). However, such methods often incur substantial computational costs and lead to increase in model sizes. Prior works also include multi-lingual LLMs optimized for several languages including Hindi : Bloom-176B (BigScienceWorkshop, 2023), Aya-23B (Aryabumi et al., 2024), Aya-101 (Üstün et al., 2024) and Aya-expanse (Dang et al., 2024). Additionally we also have several other mono-lingual and bi-lingual LLMs focused on Hindi : llama-nanda-10B (Llama-Nanda, 2024), Airavata-7B (Gala et al., 2024), (BhabhaAI, 2024), Aryabhatta-8.5B (GenVRadmin, 2024), Sarvam-2B (Sarvamai, 2024), Krutrim-2-12B (Kallappa et al., 2025) and Nemotron-mini-Hindi (Joshi et al., 2024). The key differences can be seen in Table 1.
Prior Work Model VRAM Req Approach Status
(GenVRadmin, 2024) Aryabhatta-8.5B 18 GB Vocabulary Expansion Open-Weights
(Gala et al., 2024) Airavata-7B 14 GB LoRA fine-tuning Open-Weights
Open-Dataset
(Sarvamai, 2024) Sarvam-1-2B 6 GB From Scratch Open-Weights
(Kallappa et al., 2025) Krutrim-2-12B-Instruct 50 GB Vocabulary Expansion Open-Weights
(Joshi et al., 2024) Nemotron-mini-Hindi 9 GB Continued-Pretraining Open-Weights
(Llama-Nanda, 2024) Llama-nanda-10B-chat 41 GB Block Expansion Open-Weights
Continued-Pretraining
(Dang et al., 2024) Aya-Expanse-8B 17 GB From Scratch Open-Weights
(Dang et al., 2024) Aya-Expanse-35B 66 GB From Scratch Open-Weights
Our Work placeholder-name 30 GB LoRA with modified Open-Dataset
chat template Open-Weights
Table 1: Key differences between other works regarding Hindi LLMs
@@ 3 Datasets
Despite the existence of datasets to cover several domains for Hindi (Khan et al., 2024), (Ramesh et al., 2022), we decided to experiment primarily with translated / reformatted datasets which do not prohibit usage for research/commercial purpose. This was done so that the same work can be implemented/extended to low-resource languages. Also, fine-tuning on translated data is an efficient way to adapt mPLMs to new languages, leveraging their pre-trained multilingual knowledge. (Chen and Chen, 2024). For translation, we used GPT-4o-mini (OpenAI, 2024) through Microsoft Azure  [^1] https://azure.microsoft.com/en-us/products/ai-services/openai-service/ to translate few datasets and benchmarks from English to Hindi : Big-Bench-Hard (Suzgun et al., 2022), XNLI (Conneau et al., 2018), Xl-Sum (Hasan et al., 2021). Some of the benchmarks which already have Hindi subsets were used directly : Global MMLU (Singh et al., 2024a), IndicXNLI (Aggarwal et al., 2022). Some of the publicly available datasets containing cultural and localized general knowledge like Indian legal FAQ (Aditya2411, 2024), UPSC FAQ (prnv19, 2024), IndianTAX FAQ (msinankhan1, 2024), IndianMedicines , IndiaCuisines and IndiaTravel Guide (cyberblip, 2024) were used to generate instruction-response pairs from the tabular format data using GPT-4o-mini as a part of our dataset collection. These were first translated to the other language from the original language then manually verified by multiple annotators to ensure quality in both languages. We also used a few subsets from the Aya collection (Singh et al., 2024b) i.e the translation, simplification and summarization subsets. In total the collected dataset had 3.12M samples with nearly 50:50 ratio of English and Hindi data. Around 90K samples from these cover localized and cultural knowledge. Among the rest, some domains and tasks had higher proportion in the collection. We used randomly selected subsets from those datasets while maintaining equal language ratios. After filtering the training data, we had around 485K samples of which 20% are of localized domain and cultural knowledge, while the rest are of generic tasks like math, MCQs, reasoning, summarization, rephrasing and translation.
@@ 6 Dataset Distribution and Ordering
The performance of models from initial tests didn’t vary significantly with/without being trained on math data. The performance on Math subsets of MMLU as well remained similar on both languages with/without being trained on math samples. Since we would be training on a large number of samples, we decided to still use a considerable amount of math samples. A significant performance gap was observed over boolean benchmarks with a nearly 3% increase in English and 5% increase in Hindi. Hence, we decided to use a slightly higher amount of boolean questions’ samples in the final dataset. The language ratios for each domain in the final dataset were determined based on the initial training data ratios that gave the best results. The samples of the final dataset were sorted over input lengths in ascending order with a certain number of longest samples placed in the beginning, this approach could improve batch processing efficiency and training stability (Wang et al., 2024a). This number was set equal to the total effective batch size (i.e the product of batch size and gradient accumulation steps). The samples related to local and cultural knowledge were then placed such that they are evenly spread out in the dataset except the initial batch. More info on the dataset can be found in Appendix B. The training methods and details can be found in Appendix A.
@@ 7 End Evaluation
Apart from the benchmarks seen in Table 3 and Table 4, we perform evaluations over additional benchmarks like : MMLU-Pro (Wang et al., 2024b), BigBench-Hard (Suzgun et al., 2022), MuSR (Sprague et al., 2024), GPQA (Rein et al., 2023), MATH-Hard (Hendrycks et al., 2021). We used open-llm-leaderboard  [^2] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard (Fourrier et al., 2024) for evaluation over some of the benchmarks through eval-harness framework(Gao et al., 2021). Table 8 demonstrates The performance of our models in comparison with the original models over several benchmarks. We did observe variations in the scores from open-llm-leaderboard and the corresponding benchmark scores which were self reported for the original models. We used the scores from the leaderboard for all models over those benchmarks for reproducibility a fair comparison. The evaluation methods used can be seen in Table 5.
@@
8 Generative tasks evaluation
Scarcity of genuine and authentic multilingual benchmarks of broad range of topics has been a concern for many languages. Prior works in comparison like (Llama-Nanda, 2024) have not included generative evaluations over either languages. while (Joshi et al., 2024) utilized limited generative benchmarks using LLM-as-a-judge to score the responses, with only the MT-Bench a translation task undergoing human evaluation. Further, training on translated data to test over benchmarks translated from English defeats the purpose of building multi-lingual and multi-cultural LLMs. (Aryabumi et al., 2024) too utilizes translated benchmarks for multilingual generative task evaluation, with additional human evaluation without topic/domain restriction. We have performed human-evaluation in the same way over both languages. These results can be seen in Figure 1. We performed human evaluations though third party annotators over both languages over few of the models that achieved comparably good performance over non-english discriminative tasks. A total of 3217 comparisons were done primarily in Hindi (2097) and the rest over English (1120). For a fair comparison we utilized the default hyper-parameters of each of the models.
Figure 1: Win Rates with comparable models through human evaluation.
@@ 8 Generative tasks evaluation
Our Qwen Model (14B) 90.61 94.82 88.53 90.70 75.00 87.93 52.63 36.24 44.84 64.97 25.08
Our Phi Model (14B) 97.39 92.24 87.65 87.40 75.59 88.05 52.39 39.77 49.07 66.97 23.11
@@
Our Qwen Model (14B) 74.06 81.23 84.07 78.20 53.85 74.82
Our Phi Model (14B) 81.74 89.06 86.02 78.70 56.39 78.38
@@ 9 Comparisons
For additional comparisons, we compare the performance of our models with other Hindi bilingual LLMs and other open-source LLMs which are optimized for Hindi. Due to the large variations in number of parameters of our models and other comparable models, we compare average benchmark performance versus the model size in terms of VRAM requirement. The comparisons over English and Hindi benchmarks along side our Qwen and phi models can be seen in Table 6 and Table 7. Over the benchmarks of higher difficulty, our models have consistently outperformed models over twice their size as seen in Table 6.
Benchmark Lang Qwen-2.5- Our Qwen Change Phi-4 Our Phi-4 Change
@@
Benchmark Lang Qwen-2.5- Our Qwen Change Phi-4 Our Phi-4 Change
@@ 9.1 Domain wise Performance change
The performance of our models compared to the original versions over MMLU-pro can be seen in Table 10. The type of questions the models faced through MMLU-Pro maybe of the same domain but were of different subdomains and task types compared to those in our datasets. For example, The CS benchmarks’ questions were MCQs about various areas of computer science while our training data over CS was solely from MBPP (Austin et al., 2021) which consists of a text input and a python code as an output. Further the only source of training data we used for economics consist of TAX filing FAQs over Indian context and primarily in Hindi. Hence such domains’ data usage was mentioned as N/A. The domains which had a performance boost in our models without being in training data had questions of the form of fill-mask or text completion which were similar to the training data from Winogrande-XL (Sakaguchi et al., 2021) and PIQA (Bisk et al., 2020) spanning several domains.
Figure 2: Distribution of each model’s choices over MMLU-Pro
@@ 9.2 Model biases over choices
The observations from domain wise performance changes by Phi and Qwen were significantly different. The domains which were well represented in our training data had a significant boost on both languages of MMLU. Despite training on MCQs which consist of 2-4 options, similar results of improvement were seen over MMLU-Pro which has upto 10 options. On the other hand, Phi-4 had a higher performance boost over MMLU which has the same number of options as the samples in the training data, but the performance over MMLU-Pro dropped irrespective of domain. The distribution of choices made by each of our LLMs and the corresponding original implementation can be seen in Figure 2. The instruction tuning dataset we used had an equal distribution of each of the choices among MCQ samples. The original Qwen model overwhelmingly chose from the final two options while our model was able to generalize well despite not being trained on MCQs with 10 choices. On the other hand the original phi-4 was able to perform better than its counterpart, but despite being fine-tuned with equal distribution of choices, the model displayed an inclination towards the first choice among the list of options. The extent of this bias varied between each domain significantly. More on this can be seen in Appendix D. As our phi model was fine-tuned from the original models’ instruct variants, the biases were assumed to have been carried forward. Our models were able to respond well with less biases in choices over the domains whose samples are present in large quantities in our training data. To further look into this, we tried to fine-tune the base variant of qwen-2.5-14B rather than the instruct model to see the choice distribution over MMLU-Pro, while most of our dataset’s samples of MCQs were having 4-5 samples, it was reflected in the choices made as seen in Figure 3 which demonstrates the issue within the original model similar to previous works demonstrating sensitivity on models’ sensitivity to order of choices (Pezeshkpour and Hruschka, 2024). But a well balanced instruction tuning dataset can minimize this issue or an evaluation independent of order of choices (Zheng et al., 2023). A slight tilt from left to right in Figure 2 and Figure 3 can be expected as not all questions are accompanied by 10 options with a considerable amount having less.
@@ 10 Conclusion
We demonstrate that enhancing low-resource language capabilities in LLMs is possible through targeted fine-tuning rather than complex architectural changes. Our work shows that a 12-15B parameter LLM provides an effective balance between performance and accessibility, requiring just 30GB RAM. The performance analysis reveals that our Phi-4 model excels in general-purpose tasks, while the Qwen model shows stronger adaptation to specific domains, as evidenced by the domain-wise performance changes in Table 10. Our approach of using primarily translated datasets, except for culturally specific knowledge, makes this method readily adaptable to other low-resource languages. To further push the research in low-resource languages, we release our training code, datasets, and models under commercially permissible licenses.
@@ 11 License
Our Qwen and Phi models are available through the same licenses as the models we used as a base i.e apache-2.0 and mit respectively. the models can be accessed here  [^3] Our Phi-4 model :https://huggingface.co/. The training datasets are publicly available here [^4] Datasets : https://huggingface.co/. Most datasets used for training the models have a copyleft license, with the rest having no license specified and are publicly available on huggingface.
@@
Acknowledgments
This work was supported by a prior research grant from Traversaal.ai
@@ Appendix A Model Replication
The hyper-parameters used for training can be seen below in Table 11. The initial training attempts using a portion of the data (i.e 8% samples) were done on various different devices, the final models were trained on a single H200 SXM for 55,56,54 hours each using Qwen2.5-14B-Instruct, Phi-4, Qwen2.5-14B-base respectively.
Figure 3: Distribution of response choices of our model training from Qwen-base variant over MMLU-Pro
Figure 4: Emissions : open-llm-leaderboard evaluation